Determine the ethalpy of reaction, delta H, for the formation of tungsten carbide, WC, from the elements.
W(s) + C(graphite) = WC (s)
The heats of combustion of the elements and of tungsten carbide are:
2 W(s) + 3 O2 (g) = 2 WO3 (s); delta H = -1685.8 KJ
C (graphite) + O2 (g) = CO2 (g); delta H = -393.5 KJ
2WC (s) + 5 O2 (g) = 2 WO3 (s) + 2 CO2 (g); delta H = -2391.8 KJ
Answer #1

W(s) + 3/2 O2 (g) = WO3 (s); delta H = -842.9 KJ

C (graphite) + O2 (g) = CO2 (g); delta H = -393.5 KJ

WO3 (s) + CO2 (g) = WC (s) + 5/2 O2 (g) delta H = +1195.5 kj

=====================================================================================

W(s) + C(graphite) = WC (s) delta H = -40.9 kj  

Hence  delta H of the reaction is -40.9 kj c